OpenStudy (anonymous):

2x=y+3 in slope intercept form. Please show work

OpenStudy (anonymous):

general slope intercept form is y = mx + b we are given 2x = y+3 y = 2x-3 here m = 2 b = -3

OpenStudy (anonymous):

hope you have understand :)

OpenStudy (anonymous):

Multiply both sides by -1 to get a positive y: -1(-y)= -1(-2x+3) y = 2x - 3